OpenStudy (anonymous):

Solve: 7x - 2(x + 1) = 6x + 14

OpenStudy (anonymous):

expand -2(x+1) first

OpenStudy (anonymous):

7x -2x -1 = 6x +14

OpenStudy (anonymous):

now simplify 7x-2x since they are like terms

OpenStudy (anonymous):

so that would be 7 * -2 ?

OpenStudy (anonymous):

7x-2x just like that =5x

OpenStudy (anonymous):

ok so 5x -1 = 6x + 14

OpenStudy (anonymous):

now subtract 5x from both sides what happens

OpenStudy (anonymous):

-1 = x +14

OpenStudy (anonymous):

goo can u think of what number - 14 =-1

OpenStudy (anonymous):

13

OpenStudy (anonymous):

what number + 14 = -1 i mean

OpenStudy (anonymous):

-15 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x+14=-1 you can also subtract 14 from both sides to get that result x+14-14=-1-14 x=-15

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

you're welcome